# Path Planning for a Cooperative Navigation Aid Vehicle to Assist Multiple Agents Intermittently

### v1

@@ Abstract
This paper considers the problem of planning a path for a single underwater cooperative navigation aid (CNA) vehicle to intermittently aid a set of $N$ agents to minimize average navigation uncertainty. Both the CNA and agents are modeled as constant-velocity vehicles. The agents traverse along known nominal trajectories and the CNA plans a path to sequentially intercept them. Navigation aiding is modeled by a scalar discrete time Kalman filter. During path planning, the CNA considers surfacing to reduce its own navigation uncertainty. A greedy planning algorithm is proposed that uses a heuristic based on an optimal time-to-aid, overall navigation uncertainty reduction, and transit time, to assign agents to the CNA. The approach is compared to an optimal (exhaustive enumeration) algorithm through a Monte Carlo experiment with randomized agent nominal trajectories and initial navigation uncertainty.
@@
This work has been submitted to IFAC for possible publication
@@ 1 Introduction
A cooperative navigation aid (CNA) is a surface or underwater vehicle that is equipped with a high-performance navigation system that serves as a moving reference position transponder and provides relative range and/or bearing measurements to other nearby vehicles with less capable navigation systems (Vaganay et al., 2004). Prior work on path planning for CNAs has focused on close-proximity and continuous aiding scenarios (i.e., assuming the CNA remains within communication range of the vehicles being aided at all times). The CNA-to-agent geometry during these intertasks influences the state estimate and covariance evolution onboard the receiving agents’ inertial navigation systems. Optimal steering and path planning strategies to improve localization performance during such intertasks have been explored. For example, Quenzer and Morgansen (2014) used a helming strategy guided by path inertia or the empirical observability Gramian. Chitre (2010) used approximate dynamic programming principles with a value function related to the sum-square position error of agents. The problem has also been formulated as a Markov decision process with cross-entropy learning (Teck and Chitre, 2011). Bahr et al. (2009) demonstrated a CNA path optimization strategy experimentally.
In this work, a related scenario is considered wherein the CNA can provide intermittent aiding to multiple agents. That is, we assume that there are $N$ agents requiring navigation assistance, but the CNA can only service one at a time due to their spatial separation. The goal is to design a path that minimizes average navigation uncertainty across all agents by optimizing the sequence in which all agents (or a subset of agents) are intercepted subject to a time constraint. The optimization also considers the possibility that the CNA can reduce its own navigation uncertainty (e.g., by surfacing for a GPS fix). We envision that the proposed method can be used as a high-level task allocation algorithm, to decide the sequence of agents to aid and when to surface, while existing steering strategies can optimize close-range intertasks to improve localization performance.
The contributions of this paper are: (1) a path planning problem formulation for a single CNA to intermittently aid a set of $N$ agents to minimize average navigation uncertainty with a mission time constraint, (2) a derivation of the optimal time-to-aid for a CNA aiding a single agent, and (3) a greedy algorithm that uses a novel heuristic to efficient solve the intermittent aiding problem. The performance of the proposed algorithm is evaluated through Monte Carlo simulations and compared to an optimal (exhaustive enumeration) algorithm.
The remainder of the paper is organized as follows. Section 2 describes the environment, CNA and agent motion, the navigation uncertainty models, and the cost function to be optimized. Section 3 presents intermediate results for a single CNA and agent scenario. Section 4 introduces the proposed greedy algorithm. Section 5 presents the results of numerical simulations evaluating performance. The paper is concluded in Sec. 6.
@@ 2 Problem Formulation
This work considers the problem of planning the path of a cooperative navigation aid (CNA) to aid $N$ other vehicles (referred to as agents) traveling along known nominal trajectories with known initial navigational uncertainty. In this section the environment model, vehicle motion models, navigation uncertainty models, and the problem statement are described.
@@ 2.1 Environment and CNA and Agent Motion Models
The CNA and agents are assumed to operate in the plane. The $i$th agent’s state at discrete time $k$ is the position $\bm{x}^{i}_{k}=[x^{i}_{k},y^{i}_{k}]^{\rm T}$ and its motion is described by the constant-velocity dynamics:
@@
Similarly, the CNA’s state is the position $\bm{x}^{c}_{k}=[x^{c}_{k},y^{c}_{k}]^{\rm T}$, and the CNA’s motion is described by the dynamics:
@@
where ${\bm{u}}_{k-1}^{c}=[u_{x,k-1}^{c},u_{y,k-1}^{c}]^{\rm T}$ is a velocity control input constrained to have magnitude $||{\bm{u}}_{k-1}^{c}||\leq v_{c}\Delta t$, and $v_{c}$ is the CNA’s speed. Let $\eta=v_{a}/v_{c}<1$ denote a speed ratio. The CNA’s dynamics include zero-mean Gaussian process noise ${\bm{w}}^{c}_{k}\sim\mathcal{N}({\bm{0}},{\bm{Q}}^{c})$ where ${\bm{Q}}^{c}={\bm{1}}\nu_{c}$ is a diagonal matrix with entries of variance $\nu_{c}$.
In this work, the process noise terms model the effects of navigation uncertainty (i.e., the position variables ${\bm{x}}^{c}$ and ${\bm{x}}^{i}$ represent the time-evolution of the navigation system’s estimate). The CNA is assumed to have a more capable navigation system, thus $\nu_{c}\ll\nu_{w}$.
@@ 2.2 Agent Navigation Uncertainty Model
Each agent begins at an uncertain initial position ${\bm{x}}_{0}^{i}\sim\mathcal{N}(\hat{\bm{x}}_{0|0}^{i},{\bm{P}}^{i}_{0|0})$ with mean $\hat{\bm{x}}_{0|0}^{i}=[x^{i}_{0|0}~{}~{}y^{i}_{0|0}]^{\rm T}$ and covariance ${\bm{P}}^{i}_{0|0}={\bm{1}}\nu^{i}_{0|0}$. The CNA provides absolute position measurements to the $i$th agent at time $k$:
@@
where ${\bm{H}}$ is an identity matrix, ${\bm{v}}_{k}\sim\mathcal{N}({\bm{0}},{\bm{R}}_{k})$ is a zero-mean Gaussian measurement noise vector, and ${\bm{R}}_{k}={\bm{1}}\nu_{y}+{\bm{P}}_{k|k}^{c}$. The covariance of the measurement noise, ${\bm{R}}_{k}$, consists of a constant term that depends on the relative position sensor noise $\nu_{y}$ and a time-varying term that depends on the absolute position uncertainty of the CNA itself, ${\bm{P}}_{k|k}^{c}$. This assumption models the fact that as the CNA’s own uncertainty increases so does the uncertainty in the aiding measurements it provides. Another conservative assumption is that aiding is only provided when the agent and CNA are collocated.
@@
where $\hat{\bm{x}}^{i}_{k-1|k-1}$ is the state prior, $\hat{\bm{x}}^{i}_{k|k-1}$ is the state prediction, $\hat{\bm{x}}^{i}_{k|k}$ is the state posterior, and the matrices ${\bm{F}},{\bm{G}},{\bm{Q}},{\bm{H}},{\bm{R}}$ are given in (1) and (3). If, instead, the CNA is not collocated with the $i$th agent then no measurements are transmitted. In these cases, the agent’s own navigation estimate is updated as a linear transformation of a Gaussian random vector:
@@ 2.3 CNA Navigation Uncertainty Model
$\bm{P}^{c}_{k|k}$ $=\bm{F}\bm{P}^{c}_{k-1|k-1}\bm{F}^{T}+\bm{Q}^{c}\;.$ (14)
However, the CNA is capable of periodically surfacing to obtain a GPS fix. A CNA can surface at discrete time $S$ that lasts for $M$ timesteps. For simplicity, we assume the CNA is unavailable to provide navigation updates and does not move during this time (i.e., $\bm{u}^{c}_{j}={\bm{0}}$ for $j\in\{S,S+1,...S+M\}$). After $M$ timesteps have elapsed the navigation uncertainty is re-initialized:
@@ 2.4 Problem Statement
Our goal is to design a path for the CNA to minimize average navigation uncertainty across all agents with a mission time constraint. Suppose that the CNA has prior knowledge that the agents are travelling from their nominal initial states, $\bm{\bar{x}}^{i}_{0}=\hat{\bm{x}}^{i}_{0}$, to a final state, $\bm{\bar{x}}_{T}^{i}$, along a known straight-line trajectory $\bm{\bar{x}}_{0:T}^{i}=\{\bm{\bar{x}}_{0}^{i},\bm{\bar{x}}_{1}^{i},\ldots,\bm{\bar{x}}_{T}^{i}\}$ where the final time is $T$. The CNA begins at the initial state $\bar{\bm{x}}^{c}_{0}=\hat{\bm{x}}^{c}_{0}$. The optimization problem is to determine the CNA trajectory $\bar{\bm{x}}_{0:T}^{c}=\{\bar{\bm{x}}_{0}^{c},{\bm{x}}_{1}^{c},\ldots,\bar{\bm{x}}_{T}^{c}\}$ and time to surface $S$ that
@@
To reduce complexity, we consider an optimization over a class of CNA paths that are parameterized by an ordered sequence of at most $D\leq N+1$ unique tasks. Task are from the set $\mathcal{Q}=\{0,1,2,\ldots,N\}$, where an integer task $a>0$ indicates intercepting the $i$th agent and the task $a=0$ indicates surfacing. A $D$-length task sequence is denoted by ${A}=\{a_{1},\ldots,a_{D}\}$. The set of feasible task sequences is of length $D$ or less is:
@@
where $T({A})$ denote the time to complete ${A}$. The following section describes the guidance strategy used by the CNA to complete each agent intercept task $a_{i}>0$. Thus, any task sequence ${A}\in\mathcal{F}$ can be mapped to a CNA path $\{\bar{\bm{x}}_{0}^{c},{\bm{x}}_{1}^{c},\ldots,\bar{\bm{x}}_{T}^{c}\}$ to compute the cost of (17).
@@ 3 Single-Agent Aiding Intercepts
This section presents the strategy used to intercept an agent and the change in the cost function associated with such an intertask. These results are then used in Sec. 4 to design a combined task and path planning algorithm.
@@ 3.1 Minimum Time to Intercept an Agent
$\beta_{i}={\rm acos}(\langle-{\bm{\phi}}_{i},{\bm{\theta}_{i}}\rangle)\quad\text{and}\quad\beta_{c}$ $={\rm acos}(\langle{\bm{\phi}}_{i},{\bm{\theta}_{c}}\rangle)\;,$ (18)
where ${\bm{\theta}}_{c}=[\cos\theta_{c},\sin\theta_{c},0]^{\rm T}$ is the travel direction for the CNA (to be determined). Define $\alpha=\pi-\beta_{c}-\beta_{i}$. From the sine law the intercept angle is
$\beta_{c}={\rm asin}\left(\eta{\sin\beta_{i}}\right)\;.$ (19)
@@
3.2 Agent Navigation Uncertainty after CNA Intertask
@@ 3.3 CNA Navigation Uncertainty after Surfacing
$\nu^{c}_{k|k}=\nu_{G}+k\nu_{c}\;.$ (29)
Let $S+M\leq T$ denote the time at which the vehicle is assumed to complete surfacing. Once it surfaces the variance becomes $\nu^{c}_{S+M|S+M}=\nu_{G}$. Afterwards, for $k>S+M$, the CNA navigation uncertainty is
$\nu^{c}_{k|k}=\nu_{G}+k\nu_{c}\;.$ (30)
@@ 3.4 Optimal Time-to-Aid
$J_{i}=$ $\frac{1}{T+1}\sum_{k=0}^{T}\nu^{i}_{k|k}$ (33)
$=$ $\frac{1}{T+1}\left[\sum_{k=0}^{Z-1}(\nu^{i}_{0|0}+k\nu_{w})+\sum_{k=Z}^{T}(\nu^{i}_{Z|Z}+(k-Z)\nu_{w})\right]$ (34)
$=$ $\frac{\nu_{w}T}{2}+$ (35)
@@
$\frac{\partial J_{i}}{\partial Z}$ $=\frac{2\nu_{w}Z+(\nu_{0|0}^{i}-\nu_{w}T)}{T+1}+\frac{\partial(v_{Z|Z}^{i}/\partial Z)(T-Z)-v_{Z|Z}^{i}}{T+1}$ (37)
@@ 3.5 Bounds on the Cost Function
$J_{i}^{\rm min}=J_{i}(Z_{i}^{*})$ (40)
evaluated with $\nu_{Z|Z}^{c}=\nu_{G}$. We conclude that the cost function is bounded by:
@@
[…]
@@
4 Intermittent Aiding Algorithm
The results in Section 3 established the optimal time-to-aid, $Z_{i}^{*}$, for each agent. However, reaching all of the agents at their optimal aiding times is not feasible. Instead, we propose a greedy algorithm that uses a heuristic to incorporate knowledge of $Z_{i}^{*}$ to sequentially select which agent to aid. The algorithm proceeds as follows:

### v2

# Path Planning for a Cooperative Navigation Aid Vehicle to Assist Multiple Agents Sequentially

@@ Abstract
This paper considers planning a path for a single underwater cooperative navigation aid (CNA) vehicle to sequentially aid a set of $N$ agents to minimize average navigation uncertainty. Both the CNA and agents are modeled as constant-velocity vehicles. The agents travel along known nominal trajectories and the CNA plans a path to sequentially intercept them. Navigation aiding is modeled by a scalar discrete time Kalman filter. During path planning, the CNA considers surfacing to reduce its own navigation uncertainty. A greedy planning algorithm is proposed that uses a heuristic to schedule agents to the CNA that is based on the optimal time-to-aid, the overall navigation uncertainty reduction, and the transit time. The approach is compared to an optimal (exhaustive enumeration) algorithm through a Monte Carlo experiment with randomized agent trajectories and initial navigation uncertainty.
@@
© 2024 Artur Wolek. This work has been accepted to IFAC for publication under a Creative Commons Licence CC-BY-NC-ND
@@ 1 Introduction
A cooperative navigation aid (CNA) is a surface or underwater vehicle that is equipped with a high-performance navigation system that serves as a moving reference position transponder and provides relative range and/or bearing measurements to other nearby vehicles with less capable navigation systems (Vaganay et al., 2004). Prior work on path planning for CNAs has focused on close-proximity and continuous aiding scenarios (i.e., assuming the CNA remains within communication range of the vehicles being aided at all times). The CNA-to-agent geometry during these interactions determines the state estimate and navigation uncertainty for each agent receiving the aiding. Optimal steering and path planning strategies to improve localization performance during such interactions have been explored. For example, Quenzer and Morgansen (2014) used a helming strategy guided by path inertia or the empirical observability Gramian. Chitre (2010) used approximate dynamic programming principles with a value function related to the sum-square position error of agents. The problem has also been formulated as a Markov decision process with cross-entropy learning (Teck and Chitre, 2011). Bahr et al. (2009) demonstrated a CNA path optimization strategy experimentally.
In this work, a related scenario is considered wherein the CNA can provide sequential aiding to multiple agents. That is, we assume that there are $N$ agents requiring navigation assistance, but the CNA can only service one at a time due to their spatial separation. The goal is to design a path that minimizes average navigation uncertainty across all agents by optimizing the sequence in which all agents (or a subset of agents) are intercepted subject to a time constraint. The optimization also considers the possibility that the CNA can reduce its own navigation uncertainty by surfacing for a GPS fix. We envision that the proposed method can be used as a high-level task scheduling algorithm to dectermine the sequence of agents to aid, and existing steering strategies can optimize close-range interactions to improve localization performance.
The contributions of this paper are: (1) a derivation of the optimal time-to-aid for a CNA aiding a single agent, and (2) a greedy algorithm that uses a novel heuristic to efficient solve the sequential aiding problem. The performance of the proposed algorithm is evaluated through Monte Carlo simulations.
The remainder of the paper is organized as follows. Section 2 describes the environment, CNA and agent motion, the navigation uncertainty models, and the cost function to be optimized. Section 3 presents intermediate results for a single CNA and agent scenario. Section 4 introduces the greedy algorithm. Section 5 presents the results of numerical simulations, and the paper is concluded in Sec. 6.
@@ 2 Problem Formulation
This work considers the problem of planning the path of a cooperative navigation aid (CNA) to aid $N$ other vehicles (referred to as agents) traveling along known nominal trajectories with known initial navigational uncertainty. In this section the environment model, vehicle motion models, navigation uncertainty models are described and the problem statement is introduced.
@@ 2.1 Environment and CNA and Agent Motion Models
The CNA and agents are assumed to operate in the plane, and the $i$th agent’s state at discrete time $k$ is the position $\bm{x}^{i}_{k}=[x^{i}_{k},y^{i}_{k}]^{\rm T}$. The agent’s motion is described by the constant-velocity dynamics
@@
Similarly, the CNA’s state is the position $\bm{x}^{c}_{k}=[x^{c}_{k},y^{c}_{k}]^{\rm T}$, and the CNA’s motion is described by the dynamics
@@
where ${\bm{u}}_{k-1}^{c}=[u_{x,k-1}^{c},u_{y,k-1}^{c}]^{\rm T}$ is a velocity control input constrained to have magnitude $||{\bm{u}}_{k-1}^{c}||\leq v_{c}\Delta t$, and $v_{c}$ is the CNA’s speed. Let $\eta=v_{a}/v_{c}<1$ denote a speed ratio. The CNA’s dynamics include zero-mean Gaussian process noise ${\bm{w}}^{c}_{k}\sim\mathcal{N}({\bm{0}},{\bm{Q}}^{c})$ where ${\bm{Q}}^{c}={\bm{1}}\nu_{c}$ is a diagonal matrix with entries of variance $\nu_{c}$. The process noise terms model the effects of navigation uncertainty (i.e., the position variables ${\bm{x}}^{c}$ and ${\bm{x}}^{i}$ represent the time-evolution of the navigation system’s estimate). The CNA is assumed to have a more capable navigation system, thus $\nu_{c}\ll\nu_{w}$.
@@ 2.2 Agent Navigation Uncertainty Model
Each agent begins at an uncertain initial position ${\bm{x}}_{0}^{i}\sim\mathcal{N}(\hat{\bm{x}}_{0|0}^{i},{\bm{P}}^{i}_{0|0})$ with mean $\hat{\bm{x}}_{0|0}^{i}=[x^{i}_{0|0},y^{i}_{0|0}]^{\rm T}$ and covariance ${\bm{P}}^{i}_{0|0}={\bm{1}}\nu^{i}_{0|0}$. The CNA provides absolute position measurements to the $i$th agent at time $k$:
@@
where ${\bm{H}}$ is an identity matrix, ${\bm{v}}_{k}\sim\mathcal{N}({\bm{0}},{\bm{R}}_{k})$ is zero-mean Gaussian measurement noise, and ${\bm{R}}_{k}={\bm{1}}\nu_{y}+{\bm{P}}_{k|k}^{c}$. The covariance of the measurement noise, ${\bm{R}}_{k}$, consists of a constant $\nu_{y}$ and a time-varying term that depends on the absolute position uncertainty of the CNA itself, ${\bm{P}}_{k|k}^{c}$. This assumption models the fact that as the CNA’s own uncertainty increases so does the uncertainty in the aiding measurements it provides. Aiding is assumed to occur only when the agent and CNA are collocated.
@@
where $\hat{\bm{x}}^{i}_{k-1|k-1}$ is the state prior, $\hat{\bm{x}}^{i}_{k|k-1}$ is the state prediction, $\hat{\bm{x}}^{i}_{k|k}$ is the state posterior, and the matrices ${\bm{F}},{\bm{G}},{\bm{Q}},{\bm{H}},{\bm{R}}$ are given in (1) and (3). If, instead, the CNA is not collocated with the $i$th agent then no measurements are transmitted. In these cases, the agent’s own navigation estimate is updated as:
@@ 2.3 CNA Navigation Uncertainty Model
$\bm{P}^{c}_{k|k}$ $=\bm{F}\bm{P}^{c}_{k-1|k-1}\bm{F}^{\rm T}+\bm{Q}^{c}\;.$ (14)
However, the CNA is capable of periodically surfacing to obtain a GPS fix. A CNA can surface at discrete time $S$ that lasts for $M$ timesteps. For simplicity, we assume the CNA is unavailable to provide navigation updates and does not move during this time (i.e., $\bm{u}^{c}_{j}={\bm{0}}$ for $j\in\{S,S+1,...S+M\}$). After $M$ timesteps have elapsed the navigation uncertainty is re-initialized as
@@ 2.4 Problem Statement
The goal is to design a path for the CNA to minimize average navigation uncertainty across all agents with a mission time constraint. Suppose that the CNA has prior knowledge that the agents are travelling from their nominal initial states, $\bm{\bar{x}}^{i}_{0}=\hat{\bm{x}}^{i}_{0}$, to a final state, $\bm{\bar{x}}_{T}^{i}$, along a known straight-line trajectory $\bm{\bar{x}}_{0:T}^{i}=\{\bm{\bar{x}}_{0}^{i},\bm{\bar{x}}_{1}^{i},\ldots,\bm{\bar{x}}_{T}^{i}\}$ where the final time is $T$. The CNA begins at the initial state $\bar{\bm{x}}^{c}_{0}=\hat{\bm{x}}^{c}_{0}$. The optimization problem is to determine the CNA trajectory $\bar{\bm{x}}_{0:T}^{c}=\{\bar{\bm{x}}_{0}^{c},{\bm{x}}_{1}^{c},\ldots,\bar{\bm{x}}_{T}^{c}\}$ and time to surface $S$ that
@@
To reduce complexity, the optimization is restricted to a class of CNA paths that are parameterized by an ordered sequence of at most $D\leq N+1$ unique tasks. Task are from the set $\mathcal{Q}=\{0,1,2,\ldots,N\}$, where an integer task $a>0$ indicates intercepting the $i$th agent and the task $a=0$ indicates surfacing. A $D$-length task sequence is denoted by ${A}=\{a_{1},\ldots,a_{D}\}$. The set of feasible task sequences is of length $D$ or less is:
@@
where $T({A})$ is the time to complete ${A}$. The following section describes the strategy used by the CNA to intercept each agent. Using this strategy, a sequence ${A}\in\mathcal{F}$ is mapped to a CNA path $\{\bar{\bm{x}}_{0}^{c},{\bm{x}}_{1}^{c},\ldots,\bar{\bm{x}}_{T}^{c}\}$ and cost (17).
@@ 3 Single-Agent Aiding Intercepts
This section presents the strategy used to intercept an agent and the change in the cost function associated with such an interaction. These results are then used in Sec. 4 to design a combined task and path planning algorithm.
@@ 3.1 Minimum Time to Intercept an Agent
$\beta_{i}={\rm acos}(\langle-{\bm{\phi}}_{i},{\bm{\theta}_{i}}\rangle)\quad\text{and}\quad\beta_{c}$ $={\rm acos}(\langle{\bm{\phi}}_{i},{\bm{\theta}_{c}}\rangle)\;,$
where ${\bm{\theta}}_{c}=[\cos\theta_{c},\sin\theta_{c},0]^{\rm T}$ is the travel direction for the CNA (to be determined). Note that $\alpha=\pi-\beta_{c}-\beta_{i}$. From the sine law the intercept angle is
$\beta_{c}={\rm asin}\left(\eta{\sin\beta_{i}}\right)\;.$
@@
3.2 Agent Navigation Uncertainty after CNA Interaction
@@ 3.3 CNA Navigation Uncertainty after Surfacing
$\nu^{c}_{k|k}=\nu_{G}+k\nu_{c}\;.$
$S+M\leq T$ is the time at which the vehicle is assumed to complete surfacing. Once it surfaces the variance becomes $\nu^{c}_{S+M|S+M}=\nu_{G}$. Afterwards, for $k>S+M$, the CNA navigation uncertainty is
$\nu^{c}_{k|k}=\nu_{G}+k\nu_{c}\;.$
@@ 3.4 Optimal Time-to-Aid
As described above, the evolution of the agent’s navigation uncertainty depends on when it receives a measurement $Z$. The cost (17) can be rewritten using the scalar Kalman filter notation as
@@
$J_{i}(Z)=$ $\frac{1}{T+1}\sum_{k=0}^{T}\nu^{i}_{k|k}$
$=$ $\frac{1}{T+1}\left[\sum_{k=0}^{Z-1}(\nu^{i}_{0|0}+k\nu_{w})+\sum_{k=Z}^{T}(\nu^{i}_{Z|Z}+(k-Z)\nu_{w})\right]$
$=$ $\frac{\nu_{w}T}{2}+$
@@
$\frac{\partial J_{i}}{\partial Z}$ $=\frac{2\nu_{w}Z+(\nu_{0|0}^{i}-\nu_{w}T)}{T+1}+\frac{\partial(v_{Z|Z}^{i}/\partial Z)(T-Z)-v_{Z|Z}^{i}}{T+1}$
@@ 3.5 Bounds on the Cost Function
$J_{i}^{\rm min}=J_{i}(Z_{i}^{*})$
where $J_{i}(Z_{i}^{*})$ is evaluated using (24) with $\nu_{Z|Z}^{c}=\nu_{G}$. Thus, the cost function is bounded by:
@@
It is also useful to establish the bounds of the cost function part-way along its path (i.e., at time $k$ after the CNA has completed several tasks). It is evident from the previous examples that the cost function monotonically decreases for $Z\in\{1,\ldots,Z_{i}^{*}-1\}$, reaches a minimum at $Z_{i}=Z_{i}^{*}$, and then monotonically increases for $Z\in\{Z_{i}^{*}+1,\ldots,T\}$. Thus, if the CNA has not yet aided agent $i$, then the cost $J_{i}$ at time $k$ is bounded by $J_{i}\geq J_{i}^{*}(k)$ where
@@
4 Sequential Aiding Algorithm
The results in Sec. 3 established the optimal time-to-aid, $Z_{i}^{*}$, for each agent. However, reaching all of the agents at their optimal aiding times is not feasible. Instead, we propose a greedy algorithm that uses a heuristic to incorporate knowledge of $Z_{i}^{*}$ to sequentially select which agent to aid. The algorithm proceeds as follows: